a gas sample occupies 4.20 L under a pressure of 800. torr at 77*C. At what temperature will it occupy 3.60 L at the same pressure
1
Expert's answer
2020-12-11T02:51:36-0500

V1 = 4.20 L

T1 = 77oC + 273 = 350 K

P1 = P2 = 800. torr = const

V2 = 3.60 L

T2 = unknown


Solution:

Since the pressure doesn't change, Charles' Law can be used.

Charles' Law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the final temperature, solve the equation for T2:

T2 = V2T1 / V1

T2 = (3.60 L × 350 K) / (4.20 L) = 300 K

T2 = 300 K or 27oC


Answer: a gas sample occupies 3.60 L  at 300 K (27oC).